Sol Summary – February 11th

Crew IKEA Sol Summary 11 Feb 2018

SOL 1:

Summary Title: Bringing It All Back Home

Author’s name: Atila Meszaros

Mission Status: All nominal

Sol Activity Summary:

Finally, I’m back. After a few days in the snow-capped mountains of
Colorado, you don’t have any idea how much I missed the Hab. I try not
to think about the time I have left. There is a lot of work to do and
I’m very excited about this.

The 188 crew left during the first hours of the morning, I chose my
room, I unpacked my bags and with Shannon, we planned all the
activities that we are going to carry out during this week and we went
ahead with some goals of the internship. We were working on the Ikea
project, we have already chosen several items for the kitchen and the
upper deck rooms, and we have big plans for the EVA room and the
Science Dome.

The rest of the day I’ve been working in the GreenHab. A very
difficult task. Most of the plants were completely dry, apparently
they have not been watered for at least a couple of days and others
are drowning. I will try to rescue most of them, especially the
cucumbers, which were the most damaged. As Shannon mentioned, we are
going to be keep just a few plants. Tomorrow I will be sending a
complete GreenHab Report with aquaponics status information.

Thank you very much for being with us tonight. I hope you’re having an
amazing time!


Look Ahead Plan:

Finish organizing the Green Hab, take out the Rovers and consume their
batteries to complete their cycles and finish selecting Ikea items.

Anomalies in work:


Weather: Temperature: 1°C

Crew Physical Status: Healthy

Reports to be filed:

Sol Summary
Operations Report

Support Requested: None

Operations Report – February 11th

Crew IKEA Operations Report 11 Feb 2018

SOL: 1

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control

Notes on non-nominal systems: Waiting for contractor to troubleshoot

Generator (hours run): Turned off at 9:15 am, on at 6:15 pm

Solar— SOC 92 %

Diesel – did not check %

Propane – did not check psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (trailer) – 0 gallons.

Water (static) – 400 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – no

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: none

Oil Added? n/a

ATV Fuel Used: n/a Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

Deimos rover used: No


Beginning charge:

Ending charge:

Currently charging: not sure

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging: no

Spirit rover used: No


Beginning charge:

Ending charge:

Currently charging: not sure

Opportunity rover used: No


Beginning charge:

Ending charge:

Currently charging: yes, electrical power was restored to the parking
spot on the east side of the GreenHab

Curiosity rover used: No


Beginning charge:

Ending charge:

Currently charging: Not sure

HabCar used and why, where? No

General notes and comments: Crew 188 left the Hab at 8:30 am this
morning. They did a nice job of cleaning. We are expecting rain in
the next couple of days so will make another trip for water tomorrow.

Summary of internet: nothing to report

Summary of suits and radios: nothing to report

Summary of Hab operations: The Hab is in very good condition. We
spent much of the day working on the Ikea project and have come up
with some exciting changes that will make the kitchen and EVA areas
much easier to use and maintain. We changed the air filter in the
furnace. We are expecting the new stairs, railing and upstairs
ladders to be installed this week.

Summary of GreenHab operations: Atila spent the afternoon working on
the GreenHab. Most of the large plants were dried out and wilting
while others were overwatered. The greens were well tended. We are
limiting the tomato plants to 10, pepper plants to 5 and removing the
melon plants. This is needed in order to have room for the aquaponics
system. There is a cucumber I have my eye on, and if Atila doesn’t
pick it, we will have a nice salad in a couple of days.

Summary of ScienceDome operations: There was new damage to the floor
while Crew 188 was here. There are several small pits and one larger
one in the floor.

Summary of RAM operations: Nothing to report. We are hoping the
electrician can come soon to connect the power.

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: None at this time.

Copyright © The Mars Society. All rights reserved. | Main Site